Part one:
A solid dog dish is obtained by rotating the region bounded by y=0, y=x3, x=0, x=1 about the y-axis. Determine the volume using the disk/washer method.

Part two:
Determine the volume from the last problem using the shell method.

Please write clearly. no script.. :-)
Answer #1

yax x=0 x=1 o using Disk/washer method, y0 volume, v= ((E-629) dy »v=5. j(- yat) ay » V= T. » V = T. - sv=n(1-3-o] (v = 2 an